OpenStudy (anonymous):

If the sequence is geometric, what is the common ratio? 6, 18, 54, 162, ... A. 12 B. 3 C. 27 D. not geometric

sam (.sam.):

\[\huge r=\frac{18}{6}=3\]

sam (.sam.):

\[\huge r=\frac{54}{18}=3\]

sam (.sam.):

\[\huge r=\frac{162}{54}=3\]

OpenStudy (anonymous):

A 12

OpenStudy (anonymous):

Sorry its . (3*6, = 18, 3*18 = 54, 3*54 = 162)